Donache sells cassava cakes every morning in order to raise money for tuition fee,the probability of making a sale is 0.09 for each person she approaches. Given that she approaches 18 people find the probability that she will make

1. EXACTLY SEVEN SALES

2.FIVE OR EIGHT SALES

3.NOT MORE THAN FOUR SALES


1
Expert's answer
2021-08-03T07:10:29-0400

This is a binomial distribution with n=18, p=0.09.


1) "P(X=7)=C_{18}^70.09^7(1-0.09)^{18-7}=0.0005."


2) "P(X=5)+P(X=8)=C_{18}^50.09^5(1-0.09)^{18-5}+C_{18}^80.09^8(1-0.09)^{18-8}="

"=0.0149".


3) "P(X\le4)=P(X=0)+P(X=1)+P(X=2)+P(x=3)+P(X=4)="

"=\Sigma_{i=0}^4C_{18}^i0.09^i(1-0.09)^{18-i}=0.9814."